the mean iq score for 1500 students is 90, with a standard deviation of 15. assuming the scores will have a normal curve answer the following

How many have an IQ between 75 and 105
How many have an IQ between 60 and 120
How many have an IQ over 135
1
Expert's answer
2020-03-30T11:45:55-0400

The question assumes normal distribution.

(a) Number of students with if between 75 and 105

The general normal distribution rule of 68% of units lie within one standard deviation from the mean can be used. However, we can calculate.

"Z=\frac{X-\mu}{\sigma}"

"P(75<X<105)=P(\frac{75-90}{15}<Z<\frac{105-90}{15})"

"=P(-1<Z<1)"

From z-table or =NORM.S.DIST(-1,TRUE) Excel formula, "P(Z<-1)=0.158655"

Similarly,

=NORM.S.DIST(1,TRUE) Excel formula gives "P(Z<1)=0.841345"

Thus, "P(-1<Z<1)=0.841345-0.158655=0.682697"

Thus, 68.2697% of the students lie between 75 and 105 "\frac{68.2697×1500}{100}=1024"

(b) between 60 and 120

The general normal distribution rule of 95% of units lie within one standard deviation from the mean can be used. However, we can calculate.

"P(60<X<120)=P(\frac{60-90}{15}<Z<\frac{120-90}{15})=P(-2<Z<2)"

From z table or =NORM.S.DIST(-2,TRUE) and=NORM.S.DIST(2,TRUE) Excel formulas, we have "P(Z<-2)=0.02275" and "P(Z<2)=0.97725"

Thus,"P(-2<Z<2)=0.97725-0.02275=0.9544999"

Thus, 95.44999% of students lie between 60 and 120. "\frac{95.44999×1500}{100}=1431.75" = 1432 students.

(c) over 135

"P(X>135)=1-P(Z<\frac{135-90}{15})=1-P(Z<3)"

From z table or =NORM.S.DIST(3,TRUE) Excel formula, "P(Z<3)=0.99865"

"P(X>135)=1-0.99865=0.00135"

0.135% of students have iq above 135: "\frac{0.135×1500}{100}= 2.02"

Thus, 2 students have iq above 135


Excel formulas can be used to get direct answers

(a) =(NORM.DIST(105,90,15,TRUE)-NORM.DIST(75,90,15,TRUE))×1500

(b) =(NORM.DIST(120,90,15,TRUE)-NORM.DIST(60,90,15,TRUE))×1500

(c)=(1-NORM.DIST(13,90,15,TRUE))×1500